Reports until 12:54, Tuesday 13 August 2024
GQuEST Output Filter Cavities
Torrey Cullen - posted 12:54, Tuesday 13 August 2024 - last comment - 13:44, Thursday 15 August 2024(11828)
OFC2 Realign

I scanned the laser on OFC2 for the first time since yesterday. There were higher order mode peaks on the 775 transmission PD at values higher than previously observed. This is most likely due to the earthquake that occured very near campus. I touched up the input alignment and now have almost no higher order modes seen on the TRANS PD.

Images attached to this report
Comments related to this report
Torrey Cullen - 13:44, Thursday 15 August 2024 (11835)

It's most likely 1 optic poorly clamped down that caused this. Will investigate.